Two numbers have a sum of 70, and 5 times the first number minus the second number is equal to 26. What are the two numbers?
Answer by ikleyn(52788)

.
Two numbers have a sum of 70, and 5 times the first number minus the second number
is equal to 26. What are the two numbers?
~~~~~~~~~~~~~~~

If x is the first number, then (70-x) is the second number.


Write an equation connecting the two numbers as you read the problem

    5x - (70-x) = 26     (1)


Simlify it and find x

    5x - 70 + x = 26

    5x + x = 26 + 70

       6x  =   96

        x  = 96/6 = 16.


ANSWER.  The numbers are 16 and 70-16 = 54.


CHECK.  Left side of equation (1) is  5*16 - 54 = 80 - 54 = 26.    ! Correct !

Solved.